OpenStudy (anonymous):

solve X^3-3x^2+2x-6 using factoring by grouping and the rational root theorem

OpenStudy (nurali):

x^3 -3x^2 +2x -6 = x^2(x-3) +2(x-3) ; notice the common x-3 = (x-3)(x^2+2) Hope that helps.

OpenStudy (anonymous):

thanks

OpenStudy (nurali):

Anytime.